a0poster

Version 1.21b

          Gerlinde Kettl (gerlinde.kettl@physik.uni-regensburg.de)

and

        Matthias Weiser (matthias.weiser@physik.uni-regensburg.de)^{1}

1. What is this package good for?

Printing a poster with TEX and dvips on a DIN A0 printer was quite complicated (inserting lines in the postscript file and so on). The a0poster package ants to make this easier. a0poster.cls provides fonts in sizes of 12pt (\tiny) up to 107pt (\VERYHuge). a0poster also makes sure that in math formulas the symbols appear in the right size. \textwidth and \textheight will have reasonable values. The package also creates a postscript header file for dvips which manages that the poster will be printed in the right size. Meanwhile the sizes DIN A0, DIN A1, DIN A2 and DIN A3 are also supported.

2. What the package can’t do

You still have to think about the best positioning of pictures and text.

3. System requirements and installation

Required are L^{A}TEX<1995/06/01> and dvips.

Contents of the package:

 a0poster.cls the class file

a0size.sty  adaption of the font sizes
a0 eng.tex  this manual
a0.tex  this manual in german

The files a0poster.cls and a0size.sty have to be copied in a directory that TEX scans for input files. Now the installation is done. But the package expects that dvips searches for header files in the current directory also.

4. Options

a0poster is a class like article. There are the following options:

1Thanks to Martin Eckl, Josef Rackl, Berend van der Wall and Bernhard Steininger

landscape landscape format, is default
portrait portrait format

a0b “DIN A0 big” - that’s a somewhat widened DIN A0 format that uses

the full width of an HP Designjet 650C. This is also default.

a0  DIN A0
a1  DIN A1
a2  DIN A2
a3  DIN A3

draft reduces the postscript output to DIN A4 size, so test printings can

be made with ordinary DIN A4 printers.

final  makes postscript output in original size; is default.

Therefore the beginning of the TEX file can look like that:

\documentclass[portrait,a0b,draft]{a0poster}
\usepackage{german,epsf,pstricks}
\begin{document}

The following font sizes are available:

\tiny          12pt
\scriptsize    14.4pt
\footnotesize  17.28pt
\small         20.74pt
\normalsize    24.88pt
\large         29.86pt
\Large         35.83pt
\LARGE         43pt
\huge          51.6pt
\Huge          61.92pt
\veryHuge      74.3pt
\VeryHuge      89.16pt
\VERYHuge      107pt

a0poster.cls is based on the article class, so all commands of this class can be used. Some register values were adapted to the size of the page. While TEXing a file named a0header.ps is created that is read by dvips and manages that the poster is printed in the right size.^{2}

2If this doesn’t work with the option draft, use dvips with the option -Z (if not already used in the configuration file config.ps of dvips). That also results in a smaller size of the postscript file.

If it still doesn’t work. then bigger paper sizes than DIN A3 in config.ps may be the cause. Then the draft option doesn’t work – sorry.

5. Colors, frames etc.

This package supports neither colors nor pictures, but this ist possible with other packages like pstricks of Timothy Van Zandt. With \red the following text or formula will be displayed in red. The colors red, blue, yellow, green, cyan and magenta and the grey scales white, lightgray, gray, darkgray and black are already defined, but other colors can be defined, for example

\newrgbcolor{DarkOrange}{1 .498 0}
\DarkOrange blabla

The numbers range from 0 to 1 and describe the color in the rgb system. With pstricks lots of other things can be done (frames, shadowboxes, etc.) For details, see the pstricks manual.

6. Anything else?

We are sure that this package still contains lots of bugs. So we would like to hear bug reports, complaints, suggestions etc. (just send an email to

gerlinde.kettl@physik.uni-regensburg.de).

7. History

Version 1.21b:

• Draft option available in combination with option a3.

• Modified ghostview removed.

   • a0size.tex renamed to a0size.sty.

Version 1.2b:

• Support of DIN A0, DIN A1, DIN A2 und DIN A3.

• Option draft available.

• Modified ghostview for viewing the poster.